# Time Evolution of Multi-Party Entanglement Signals

### v1

@@ 1 Introduction
The evolution of the triple information in time-dependent holographic states dual to asymptotically AdS black holes was previously considered in Balasubramanian:2011at . [^3] In the precise setup of Balasubramanian:2011at the membrane theory does not apply because they do not examine a late time, large region limit. For the analysis there it was sufficient to use the HRT formula Hubeny:2007xt for entanglement entropy of intervals in holographic theories because the triple information is constructed from such quantities. To construct the entanglement signals studied in Balasubramanian:2024ysu we need to additionally compute reflected entropies Dutta:2019gen ; Akers:2019gcv ; Hayden:2021gno ; Gadde:2023zzj ; Harper:2024ker ; Iizuka:2025ioc . The time evolution of reflected entropy was first studied in 2d CFT in Kudler-Flam:2020url . More recently, the membrane theory framework was enriched to enable computation of reflected entropy Jiang:2024tdj . We will use this method to study the dynamics of the multiparty signals considered in Balasubramanian:2024ysu .
@@ 2 Review of multi-party entanglement signals
where $S_{i}$ is the permutation-invariant combination of all the entropies on sets of $i$ parties chosen from the $n$ parties. [^6] Note that it has been recently shown that of these quantities, only the mutual information is monotonic until the partial trace Hernandez-Cuenca:2023iqh . On pure states $I_{n}=2I_{n-1}$, where $I_{n-1}$ is evaluated on the reduced density matrix obtained by tracing out any one party. For $n$ odd $I_{n}=0$ on pure states, so the $n$-information is not helpful. For odd $n$ we can define instead the $n$-residual information, by tracing out one party and taking the difference between $I_{n-1}$ evaluated on the canonical purification and $I_{n-1}$ on the density matrix,
@@ 4.1 A single finite region
So we have [^11] In the case of adjacent intervals, $S_{R}$ is not UV finite Dutta:2019gen ; Kudler-Flam:2020url . In this case, we can regularize the EWCS the same way as (H)RT, by subtracting UV divergences of the form (3.6). These UV divergences cancel in $R_{3}$.
@@ 4.2 Two finite regions
The $AB$ entanglement wedge is always connected. The cross-section is initially a vertical surface hanging down from the boundary. It transitions to a surface extending across the $B$ region when $t=\ell_{B}$, see Figure 9. When $S(AB)$ saturates at $t=\ell_{A}+\ell_{B}$, the cross-section goes back to hanging down from the boundary but just to the $v_{B}$ cone membrane corresponding to $S(AB)$ (see Figure 10). However, this cross-section is no longer described by membrane theory, as it does not grow with $t$ or $x$ in the scaling limit. Explicit calculations in the holographic case in Appendix A show that we have $S_{R}=0$ in the scaling limit in this case (after subtracting the UV divergences as discussed in footnote 11), so
@@ 8 Discussion
Since, in the absence of initial (bipartite or multipartite) entanglement, the quasiparticle picture leads to purely bipartite entanglement while, as we showed, the membrane picture often produces non-trivial multipartite entanglement, we can regard the generation of multiparty entanglement as a common feature of scrambling dynamics. It would be interesting to investigate cases with initial multipartite entanglement in the quasiparticle Casini:2015zua and membrane pictures in future works.
@@ Appendix A Saturation of $S_{R}$ in finite adjacent intervals
where $\lambda\in(-\infty,+\infty)$. To find the minimal cross-section bisecting the $AB$ entanglement wedge, we minimise the geodesic distance between $K$  [^23] To regularize the UV divergence, we consider the point $K$ at $z=\epsilon\ll 1$ in the asymptotic region. and an arbitrary point $G$ on the parametric curve (A.3). The geodesic distance formula in planar BTZ black brane is Mezei:2016zxg ; Shenker:2013pqa
@@
Now, let us consider saturated $S_{R}$ in the adjacent interval case in $d>2$ with strip entangling subregions. In the large region limit, the intersection point $z_{0}$ between the EWCS and the RT surface is close to the horizon, $z_{0}=1-\delta$ ($\delta\ll 1$). Following the intuition above, the EWCS is also almost radial in the scaling limit when $d>2$. Thus, it would be convenient to parametrize the RT surface by $x(z)$. The area functional is given by
@@
To sum up, we find that in $d\geq 2$ in the large region limit, the regularized area of the saturated EWCS is 0. As these EWCS as shown in Figure 10 (top) are radial and reach close to the horizon, their projections to the boundary along constant infalling time are vertical lines dangling from $K$, see Figure 10 (bottom). This is due to the near-horizon logarithmic divergence of infalling time $v$ (A.2Jiang:2024tdj . Nevertheless, these cross-sections are no longer described by membrane theory, as they do not grow with $t$ or $x$ in the scaling limit. In particular, these vertical “membranes” should not be confused with the vertical membranes with tension $\mathcal{E}(0)=v_{E}$ Mezei:2018jco obtained from projections of the HRT surfaces in the black hole interior Hartman:2013qma ; Liu:2013iza ; Liu:2013qca . The latter reaches the $t=0$ slice in membrane theory, computing entanglement entropy $S=s_{\rm th}v_{E}t$ that grows linearly with time.

### v2

@@ 1 Introduction
The evolution of the triple information in time-dependent holographic states dual to asymptotically AdS black holes was previously considered in Balasubramanian:2011at . [^3] In the precise setup of Balasubramanian:2011at the membrane theory does not apply because they do not examine a late time, large region limit., [^4] See also for $n$ For the analysis there it was sufficient to use the HRT formula Hubeny:2007xt for entanglement entropy of intervals in holographic theories because the triple information is constructed from such quantities. To construct the entanglement signals studied in Balasubramanian:2024ysu we need to additionally compute reflected entropies Dutta:2019gen ; Akers:2019gcv ; Hayden:2021gno ; Gadde:2023zzj ; Harper:2024ker ; Iizuka:2025ioc . The time evolution of reflected entropy was first studied in 2d CFT in Kudler-Flam:2020url . More recently, the membrane theory framework was enriched to enable computation of reflected entropy Jiang:2024tdj . We will use this method to study the dynamics of the multiparty signals considered in Balasubramanian:2024ysu .
@@ 2 Review of multi-party entanglement signals
where $S_{i}$ is the permutation-invariant combination of all the entropies on sets of $i$ parties chosen from the $n$ parties. [^7] Note that it has been recently shown that of these quantities, only the mutual information is monotonic until the partial trace Hernandez-Cuenca:2023iqh . See also Ju:2024hba ; Ju:2025tgg for further studies on $n$-information. On pure states $I_{n}=2I_{n-1}$, where $I_{n-1}$ is evaluated on the reduced density matrix obtained by tracing out any one party. For $n$ odd $I_{n}=0$ on pure states, so the $n$-information is not helpful. For odd $n$ we can define instead the $n$-residual information, by tracing out one party and taking the difference between $I_{n-1}$ evaluated on the canonical purification and $I_{n-1}$ on the density matrix,
@@ 4.1 A single finite region
So we have [^12] In the case of adjacent intervals, $S_{R}$ is not UV finite Dutta:2019gen ; Kudler-Flam:2020url . In this case, we can regularise the EWCS the same way as (H)RT, by subtracting UV divergences of the form (3.6). These UV divergences cancel in $R_{3}$.
@@ 4.2 Two finite regions
The $AB$ entanglement wedge is always connected. The cross-section is initially a vertical surface hanging down from the boundary. It transitions to a surface extending across the $B$ region when $t=\ell_{B}$, see Figure 9. When $S(AB)$ saturates at $t=\ell_{A}+\ell_{B}$, the cross-section goes back to hanging down from the boundary but just to the $v_{B}$ cone membrane corresponding to $S(AB)$ (see Figure 10). However, this cross-section is no longer described by membrane theory. Intuitively, this is because it does not grow with $t$ or $x$ in the scaling limit. In Appendix A, we show using holography that for this cross-section $S_{R}=0$ in the scaling limit (after subtracting the UV divergences as discussed in footnote 12). So
@@ 8 Discussion
Since, in the absence of initial (bipartite or multipartite) entanglement, the quasiparticle picture leads to purely bipartite entanglement while, as we showed, the membrane picture often produces non-trivial multipartite entanglement, we can regard the generation of multiparty entanglement as a common feature of scrambling dynamics. It would be interesting to investigate cases with initial multipartite entanglement in the quasiparticle Casini:2015zua ; Berthiere:2024sio and membrane pictures in future works.
@@ Appendix A Saturation of $S_{R}$ in finite adjacent intervals
where $\lambda\in(-\infty,+\infty)$. To find the minimal cross-section bisecting the $AB$ entanglement wedge, we minimise the geodesic distance between $K$  [^24] To regularise the UV divergence, we consider the point $K$ at $z=\epsilon\ll 1$ in the asymptotic region. and an arbitrary point $G$ on the parametric curve (A.3). The geodesic distance formula in planar BTZ black brane is Mezei:2016zxg ; Shenker:2013pqa
@@
Now, let us consider saturated $S_{R}$ in the adjacent interval case in $d>2$ with strip entangling subregions. In the large region limit, the intersection point $z_{0}$ between the EWCS and the RT surface is close to the horizon, $z_{0}=1-\delta$ ($\delta\ll 1$). Following the intuition above, the EWCS is also almost radial in the scaling limit when $d>2$. Thus, it would be convenient to parametrise the RT surface by $x(z)$ (instead of $z(x)$). The area functional is given by
@@
To sum up, we find that in $d\geq 2$ in the large region limit, the regularised area of the saturated EWCS is 0. As these EWCS as shown in Figure 10 (top) are radial and reach close to the horizon, their projections to the boundary along constant infalling time are vertical lines dangling from $K$, see Figure 10 (bottom). This is due to the near-horizon logarithmic divergence of infalling time $v$ (A.2Jiang:2024tdj . Nevertheless, these cross-sections are no longer described by membrane theory for holographic CFT Mezei:2018jco ; Jiang:2024tdj . This is because to derive the membrane tension (3.1) from the area of the HRT surface in the scaling limit, one needs to integrate out the radial bulk $z$ degree of freedom. However, from e.g. (A.9) above, it is clear that for the almost-radial cross-section shown in Figure 10, the radial bulk $z$ degree of freedom cannot be integrated out. The reason why this cross-section fails to be captured by membrane theory in other chaotic systems than holographic CFTs is technically less clear. Intuitively, one can understand the reason as cross-sections of this kind do not grow with $t$ or $x$ in the scaling limit. Hence, these vertical “membranes” should not be confused with the vertical membranes with tension $\mathcal{E}(0)=v_{E}$ Mezei:2018jco obtained from projections of the HRT surfaces in the black hole interior Hartman:2013qma ; Liu:2013iza ; Liu:2013qca . The latter reaches the $t=0$ slice in membrane theory, computing entanglement entropy $S=s_{\rm th}v_{E}t$ that grows linearly with time.